OpenStudy (ujjwal):

A mass m hangs with the help of a string wrapped around a pulley on a frictionless bearing. The pulley has mass 'm' and radius 'R' . Assuming pulley to be a perfect circular disc, the acceleration of mass if the string doesn't slip on the pulley is a)g/3 b)3g/2 c)g d)2g/3

mathslover (mathslover):

K just wait

mathslover (mathslover):

Should i upload the direct answer or explain u step by step ?

mathslover (mathslover):

isn't this of AIEEE 2011 ?

OpenStudy (ujjwal):

explain!

mathslover (mathslover):

\[\large{\frac{mg}{m+\frac{1}{R^2}}=\frac{mg}{m+\frac{m}{2}}=\frac{2g}{3}}\]

mathslover (mathslover):

d) 2g/3 is the correct answer

OpenStudy (ujjwal):

Ok. How do you get it? I didn't understand!

OpenStudy (anonymous):

\[\frac{I}{R^2}\]

mathslover (mathslover):

I dont want to cheat any1 ; http://brilliantiitjee.com/docs/aieee-solution-2011.pdf if it does not help then u can ask me

OpenStudy (ujjwal):

I still don't get it. Can you explain a bit?

mathslover (mathslover):

Do you know that mg-T=ma ?

OpenStudy (ujjwal):

Ok, tell me what is \[\frac{1}{2}mR^2 \alpha\]?? To be precise, tell me what is \(\alpha\) ??

mathslover (mathslover):

Awesome question that is what i am thinking about wait

OpenStudy (ujjwal):

Ok!

mathslover (mathslover):

alpha is angular acceleration

mathslover (mathslover):

\[\large{\frac{1}{2}mR^2\alpha}\]is related to momentum of inertia

mathslover (mathslover):

got it now ?

OpenStudy (ujjwal):

I had read moment of inertia a year back. An i didn't enjoy it much! so, I have forgotten almost everything related to moment of inertia. :(

mathslover (mathslover):

no problem search it on internet u will get everything there

OpenStudy (ujjwal):

Thanks!

OpenStudy (ujjwal):

Is \[I=\frac{1}{2}mR^2\]?? Is it the moment of inertia of disc?

mathslover (mathslover):

yes

OpenStudy (ujjwal):

How do they get it? And isn't moment of inertia = mR^2 ?? Where does (1/2) come from?

OpenStudy (anonymous):

\[I=\int\limits r^2 dm\] for a thin disc \[I= \int\limits_{0}^{R} \int\limits_{0}^{2\pi} r^2 \rho \ r \ dr \ d \theta= \frac{R^4}{4} 2 \pi \rho= \frac{1}{2} M R^2 \\ since \\ M=\rho 2 \pi R^2 \]

OpenStudy (anonymous):

sorry M=rho pi R^2 *******

OpenStudy (vincent-lyon.fr):

\(\huge \tau\large=I\:\alpha\) is the equivalent of \(\Large F\large=m\:a\) for rotating bodies. I think you have to memorise the moment of inertia of a solid disc/cylinder rotating about its axis: \(\Large I=\frac12\large mR^2\)

OpenStudy (anonymous):

so the T in mg - T = ma is torque and not tension?

OpenStudy (ujjwal):

Its tension! TR is the torque.

OpenStudy (anonymous):

i thought that was tension*radius! i was wondering why we're taking that!

OpenStudy (ujjwal):

Tension \(\times\) radius= torque.

OpenStudy (anonymous):

oh yes! T = r x F....right! now i get it. thanks!

OpenStudy (ujjwal):

:)